6x+3y=-3
-5x+y=20
6x+3y=-3, -5x+y=20
add 5 to both ends of 2nd equation
-5x+5x+y=20+5x
y = 20+5x
substitute y in the first equation
6x+3(20+5x)=-3
6x+60+15x=-3
21x=-63
divide by 21
x= -63/21
x= -3
plug value of x into the second equation
-5(-3)+y=20
15+y=20
y=5
Solve by substitution method

Found 2 solutions by mananth, needhelpfastplease:
Answer by mananth(16946)

6x+3y=-3
...
-5x+y=20
+5x
-5x+5x+y=20+5x
y = 20+5x
..
substitute y in first equation
6x+3(20+5x)=-3
6x+60+15x=-3
21x=-63
/21
x= -63/21
x= -3
...
plug value of x in second equation
5x+y=20
5(-3)+y=20
-15+y=20
+15
-15+15+y = 20+15
y=35
....
m.ananthe@hotmail.ca

..

Answer by needhelpfastplease(1)

6x+3y=-3, -5x+y=20
add 5 to both ends of 2nd equation
-5x+5x+y=20+5x
y = 20+5x
substitute y in the first equation
6x+3(20+5x)=-3
6x+60+15x=-3
21x=-63
divide by 21
x= -63/21
x= -3
plug value of x into the second equation
-5(-3)+y=20
15+y=20
y=5
CHECK.... 6*-3+3*5= -3 (remember pemdas so it looks like -18+15=-3)
......... -5*-3+5=20 (15+5=20)
ANSWER x=-3 Y=5